# Low-Rank Gradient Compression with Error Feedback for MIMO Wireless Federated Learning

### v1

@@ III-E Ota-LC: Signal Design
$N_{cu}=\frac{(m+n)r}{2N_{t}},$ (16)
where all antennas transmit different entries of each ${\mathbf{C}}_{k}^{(t)}$.
Let ${\mathbf{A}}^{(t)}\in\mathbb{C}^{N_{r}\times N_{t}}$ denote the receive beamforming matrix and let ${\mathbf{B}}_{k}^{(t)}\in\mathbb{C}^{N_{t}\times N_{t}}$ denote the transmit beamforming matrix of device $k$ at iteration $t$. We assume that the server can accurately estimate the CSI of each device to obtain the beamforming matrices. We use the over-the-air beamforming method in [24], and hence the beamforming matrices are designed as
${\mathbf{A}}^{(t)}$ $=\sqrt{\max_{k}\frac{1}{P_{0}}{\rm tr}\left((({{\mathbf{F}}^{(t)}})^{\dagger}{\mathbf{H}}_{k}^{(t)}[{\mathbf{H}}_{k}^{(t)}]^{\dagger}{\mathbf{F}}^{(t)})^{-1}\right)}{\mathbf{F}}^{(t)},$ (17)
@@
where ${\mathbf{F}}^{(t)}\in\mathbb{C}^{N_{r}\times N_{t}}$ is given as ${\mathbf{F}}^{(t)}=\sum_{k=1}^{K}\lambda_{\min}(\Sigma_{k}^{2}){\mathbf{U}}_{k}{\mathbf{U}}_{k}^{\dagger}$, where ${\mathbf{U}}_{k}$ and $\Sigma_{k}$ are obtained by compact singular value decomposition (SVD) of ${\mathbf{H}}_{k}$ as ${\mathbf{H}}_{k}={\mathbf{U}}_{k}\Sigma_{k}{\mathbf{V}}_{k}^{\dagger}$. Then, the transmit signal of each device at each iteration $t$ is
${\mathbf{X}}_{k}^{(t)}={\mathbf{B}}_{k}^{(t)}{{\mathbf{S}}}_{k}^{(t)}\in\mathbb{C}^{N_{t}\times N_{cu}},$ (18)
where ${{\mathbf{S}}}_{k}^{(t)}\in\mathbb{C}^{N_{t}\times N_{cu}}$ is obtained by reshaping the matrix ${\mathbf{C}}_{k}^{(t)}$. We assume the inequality, $N_{cu}\leq\tau$, which can be satisfied by choosing suitably the rank $r$.
@@ IV Numerical Experiments
We consider a wireless federated learning system with one server and $K=20$ edge devices. In each communication round, only a randomly selected half of the devices upload their local gradients to the server. The transmit SNR $\rho={P_{0}}/{N_{0}}$ is set to $20$ dB. The number of transmitter antennas of each device is set as $N_{t}=8$, and the number of the receive antennas of the server as $N_{r}=80$. The entries of the communication channels ${\mathbf{H}}_{k}^{(t)}$ are generated from i.i.d. $\mathcal{CN}(0,1)$ for all $k$ and $t$. The channels are assumed to be constant within each iteration $t$.
The experiments were performed on the classification task for the Cifar-10 dataset. We flipped and cropped the dataset for data augmentation to avoid overfitting as in [26, 27]. We use the CNN model ResNet18, setting the learning rate as $\eta=0.1$, and the weight $\beta^{(t)}$ in LC update to $0.5$ for all $t$. We control the number of parameters uploaded by the local devices by changing the value of the rank $r$. Our method is compared with SGD without compression, as well as with three existing gradient compression methods for wireless federated learning, as detailed next.
@@
$\bullet$ BLUE-CS [8]: Similar to Ota-CS, BLUE-CS utilizes Turbo-CS for gradient compression, but, during data transmission, it adopts the best linear unbiased estimator (BLUE) algorithm to independently recover transmitted signals at the server, rather than leveraging over-the-air computing.
@@
We first provide a convergence performance comparison by evaluating training loss and test accuracy against the number of communication rounds $T$ for $N_{cu}=9082$ channel uses and $r=20$. It can be observed in Fig. 2 that the proposed Ota-LC approach has a similar test accuracy at the last round as SGD while demonstrating a faster convergence rate than other methods. For instance, when aiming a learning accuracy of $80\%$ in Fig. 3, $300$ iterations are needed for BLUE-CS and Ota-CS, $165$ iterations for Ota-RLC, while Ota-LC requires only $115$ iterations. Thus, to attain equivalent performance, Ota-CS reduces the communication overhead by $30\%$.
Figure 2: Training loss (a) and test accuracy (b) as a function of the number of communication rounds $T$ ($N_{cu}=9082$, $r=20$).
Then, we compare Ota-LC with the mentioned benchmarks for different numbers of channel uses $N_{cu}$ over $T=150$ communication rounds. The results in Fig. 3 show that Ota-LC performs better under the same communication overhead, especially in the regime of a low number of communication resources $N_{cu}$, which corresponds to the case of a more aggressive compression ratio.
Figure 3: Test accuracy versus the number of channel uses $N_{cu}$ ($T=150$).
@@ V Conclusion
In this letter, we have proposed Ota-LC, a gradient compression and transmission method that reduces the communication overhead for MIMO wireless federated learning. Experimental results show Ota-LC converges faster and has better performance than existing benchmarks [8, 7, 13], while having smaller communication and computational costs. Future work may investigate the fully decentralized extension of Ota-LC.

### v2

@@ III-B1 Signal Processing
$N_{cu}={(m+n)r}/{(2N_{t})},$ (18)
where all antennas transmit different entries of each ${\mathbf{C}}_{k}^{(t)}$. Then the matrix ${{\mathbf{C}}}_{k}^{(t)}$ in (17) is reshaped into ${\mathbf{S}}_{k}^{(t)}$ as
${\mathbf{S}}_{k}^{(t)}=[{\mathbf{c}}_{k,1}^{(t)},\ldots,{\mathbf{c}}_{k,N_{cu}}^{(t)}]\in\mathbb{C}^{N_{t}\times N_{cu}},$ (19)
where ${\mathbf{c}}_{k,b}^{(t)}={\mathbf{c}}_{k}^{(t)}((b-1)N_{t}+1:bN_{t})\in\mathbb{C}^{N_{t}},b=1,\ldots,N_{cu}$, and ${\mathbf{c}}_{k}^{(t)}\in\mathbb{C}^{N_{t}N_{cu}}$ is the obtained by vectorizing ${\mathbf{C}}_{k}^{(t)}$.
Let ${\mathbf{A}}^{(t)}\in\mathbb{C}^{N_{r}\times N_{t}}$ denote the receive beamforming matrix and let ${\mathbf{B}}_{k}^{(t)}\in\mathbb{C}^{N_{t}\times N_{t}}$ denote the transmit beamforming matrix of device $k$ at iteration $t$. We assume that the server can accurately estimate the CSI of each device to obtain the beamforming matrices. We use the over-the-air beamforming method in [22]. The devices apply zero-forcing beamforming to transmit the local factorization matrices, while the server applies linear mean-squared error beamforming, yielding the beamforming matrices
${\mathbf{A}}^{(t)}$ $=\sqrt{\max_{k}\frac{1}{P_{0}}{\rm tr}\left(({{\mathbf{F}}^{(t)}})^{\dagger}{\mathbf{H}}_{k}^{(t)}[{\mathbf{H}}_{k}^{(t)}]^{\dagger}{\mathbf{F}}^{(t)})^{-1}\right)}{\mathbf{F}}^{(t)},$ (20)
@@
where ${\mathbf{F}}^{(t)}=\sum_{k=1}^{K}\lambda_{\min}(\Sigma_{k}^{2}){\mathbf{U}}_{k}{\mathbf{U}}_{k}^{\dagger}\in\mathbb{C}^{N_{r}\times N_{t}}$, where ${\mathbf{U}}_{k}$ and $\Sigma_{k}$ are obtained by compact singular value decomposition (SVD) of ${\mathbf{H}}_{k}$ as ${\mathbf{H}}_{k}={\mathbf{U}}_{k}\Sigma_{k}{\mathbf{V}}_{k}^{\dagger}$. Then, the transmit signal of each device at each iteration $t$ is
${\mathbf{X}}_{k}^{(t)}={\mathbf{B}}_{k}^{(t)}{{\mathbf{S}}}_{k}^{(t)}\in\mathbb{C}^{N_{t}\times N_{cu}}.$ (21)
We assume the inequality, $N_{cu}\leq\tau$, which can be satisfied by choosing suitably the rank $r$.
@@ III-B2 Error feedback
By plugging (23) into (24), one can see that error feedback attempts to compensate for compression errors, communication distortion, and client drifts, which are accumulated over the iterations.
@@ IV Numerical Experiments
We consider a wireless federated learning system with one server and $K=10$ edge devices. In each communication round, only a randomly selected half of the devices upload their local gradients to the server. The transmit SNR $\rho={P_{0}}/{N_{0}}$ is set to $20$ dB. The number of transmitter antennas of each device is set as $N_{t}=8$, and the number of the receive antennas of the server as $N_{r}=8$. The entries of the communication channels ${\mathbf{H}}_{k}^{(t)}$ are generated from i.i.d. $\mathcal{CN}(0,1)$ for all $k$ and $t$. The channels are assumed to be constant within each iteration $t$. The experiments were run with PyTorch on NVIDIA 3090 GPUs and the code can be found at https://github.com/MingzhaoGuo/Ota-LC.
The experiments are performed on the classification task for the Cifar-10 and Cifar-100 datasets in the i.i.d. and non-i.i.d. cases respectively. In the i.i.d. experiment, the data are randomly assigned to each user. In the non-i.i.d. experiment, the local datasets are divided using the Dirichlet distribution with scaling parameter $0.9$ [23]. We use the ResNet18 model, setting the learning rate as $\eta=0.1$ for the Cifar-10 dataset and $\eta=0.005$ for the Cifar-100 dataset, and the weight $\beta^{(t)}$ in LC update to $0.5$ for all $t$. We control the number of parameters uploaded by local devices by changing the value of the rank $r$. Our method is compared with SGD without compression, as well as with five existing gradient compression methods with error feedback for wireless federated learning, as detailed next.
@@
$\bullet$ PowerSGD[14]: PowerSGD also compresses the local gradient by finding a low-rank approximation. Unlike Ota-LC, at each iteration, PowerSGD utilizes one step of power iteration, requiring two rounds of transmission and aggregation. The local computation has a complexity of order $\mathcal{O}(nmr)$, and PowerSGD leverages digital communications.
$\bullet$ Top-$K$[7] and Rand-$K$[8]: The local computational cost of Top-$K$ has an order of $\mathcal{O}(mn\log(mn))$ due to the sorting, and Randk has an order of $\mathcal{O}((m+n)r)$. These methods leverage digital communications.
Figure 2: Test accuracy as a function of the number of $N_{cu,total}$ ($r=20$).
We first provide a convergence performance comparison by evaluating training loss and test accuracy against the total channel uses $N_{cu,total}=TN_{cu}$ and $r=20$. It can be observed in Fig. 2 that Ota-LC demonstrates a faster convergence rate than other methods in both i.i.d. and non-i.i.d. scenarios. For instance, when aiming a learning accuracy of $70\%$ in Fig. 2b, $1.5\times 10^{7}$ channel uses are required for Ota-RLC and PowerSGD, while Ota-LC requires only $1\times 10^{7}$ channel uses, reducing the communication overhead by $33\%$.
Figure 3: Test accuracy versus the number of $N_{cu}$ (a), and SNR (b) ($T=250$).
Then, we compare Ota-LC with the mentioned benchmarks for different numbers of channel uses $N_{cu}$ over $T=250$ and SNR $\rho=20$ dB for Cifar-100 under the i.i.d. scenario. The results in Fig. 3a show that Ota-LC performs better under the same communication overhead, especially in the regime of a low number of communication resources $N_{cu}$, and that error feedback helps improve performance. As $N_{cu}$ increases, the performance of Ota-LC gradually improves, approaching the performance of uncompressed gradient (SGD). We also compare Ota-LC with the benchmarks for different SNR under $T=250$ and $N_{cu}=4000$. Fig. 3b shows that Ota-LC outperforms other algorithms even in low SNR regimes.
@@ V Conclusion
In this letter, we have proposed Ota-LC, a gradient compression and transmission method that reduces the communication overhead for MIMO wireless federated learning. Experimental results show that Ota-LC converges faster and has better performance than existing benchmarks while having smaller communication and computational costs. Future work may investigate the fully decentralized extension of Ota-LC and consider the effect of error feedback strength to enhance performance.